OpenStudy (anonymous):

Find the equation of the line that passes through the points (-4,1) and (-2,-5)

OpenStudy (anonymous):

You can find the gradient of the line by y2-y1/x2-x1 Then substitute into the general equation y=mx+c Substitute y,x, and m to find c

OpenStudy (anonymous):

point slope form y - y1 = m(x - x1) where m = slope of line

OpenStudy (anonymous):

y1 = 1, x1 = -4